If A firm with market power has estimated the following demand function for its product:
Q = 12,000 – 4,000 P where P = price per unit and Q = quantity demanded per year.
The firm’s total costs are $4,000 when nothing is being produced. These costs increase by 50 cents for each unit produced. What is the equation for the firms total cost function, marginal cost function, total revenue and marginal revenue (in terms of Q)
1
Expert's answer
2012-10-11T08:24:08-0400
Total revenue can be expressed by multiplying Quantity and Price
Total Revenue = Q*P= 12,000P-4,000P2
Marginal revenue function is derivative of total cost function:
Marginal revenue = 12,000-8,000P
Total cost = 4,000+0.5Q
Marginal cost is derivative of total cost
Marginal cost = 0.5